# OceanAI: A Conversational Platform for Accurate, Transparent, Near-Real‑Time Oceanographic Insights

### v1

@@ 1 Introduction
The ocean plays a pivotal role in regulating Earth’s climate, supporting biodiversity, and sustaining coastal economies. Accurate understanding of oceanographic processes underpins climate forecasting, hazard preparedness, and sustainable resource management (Hurrell et al., 2009; Roemmich et al., 2019). Over the past two decades, advances in satellite remote sensing, in-situ observation networks (e.g., Argo (Wong et al., 2020)), and reanalysis programs(e.g., ERA5 (Hersbach et al., 2020) and GLORYS (European Union-Copernicus Marine Service, 2018)) have dramatically expanded the availability of high-resolution ocean data. However, these datasets remain technically challenging to access and work with due to their complex formats and spatiotemporal structure, creating barriers for non-expert users and overhead even for domain experts.
Existing tools for accessing oceanographic information fall broadly into three categories:general-purpose large language models (LLMs), domain-specific data servers, and hybrid geospatial analysis environments. General-purpose large language models (LLMs), such as GPT-4 (OpenAI et al., 2024) and LLaMA (Touvron et al., 2023), offer intuitive natural language interaction but lack direct access to authoritative datasets. Without integration to structured data sources, these models cannot reliably execute parameterized scientific queries, and their outputs are prone to hallucination Ji et al. (2023), producing unverifiable or incorrect statements. Domain-specific platforms, such as the NOAA THREDDS Data Server [^1] https://www.ncei.noaa.gov/thredds/catalog/catalog.html and the NOAA Open Data Dissemination Program [^2] https://www.noaa.gov/information-technology/open-data-dissemination, provide structured, transparent access to observational and model data but require substantial technical expertise to navigate file hierarchies, parse specialized formats (e.g., NetCDF, GRIB), and perform downstream analysis. Hybrid geospatial analysis environments, such as Google Earth Engine (Gorelick et al., 2017), provide access to global geospatial datasets—including some oceanographic layers—and support large-scale data processing and interactive visualization. However, they typically require users to write JavaScript or Python code and do not offer conversational interfaces or direct parameterized querying of authoritative ocean datasets.
@@ 2 Ocean Background
Sea surface temperature (SST) is widely used in earth sciences from ocean ecosystem conservation to weather forecasting. SST is usually measured by satellite sensors as well as in-situ platforms such as buoys and drifters. Prominent sources include NOAA’s Coral Reef Watch and the Group for High Resolution Sea Surface Temperature(GHRSST), which offer daily global SST data at 5 km and finer resolution in NetCDF format. These files include spatial/temporal coordinates, quality flags, and uncertainty estimates, requiring coding expertise with scientific data packages such as xarray or netCDF4 for access (NOAA Coral Reef Watch, 2023).
@@
  • El Niño-Southern Oscillation (ENSO): Characterized by prolonged SST anomalies in the eastern tropical Pacific, ENSO events influence global atmospheric circulation and provoke extreme weather like droughts in South Asia and flooding in South America (McPhaden et al., 2006; Rajeevan and Pai, 2007).
  • Monsoon variability: Warmer SSTs in the Indian and Pacific Oceans enhance monsoon rainfall; cooler SSTs can suppress or delay monsoon onset, hindering agricultural production and affecting large populations (Rajeevan and Pai, 2007; Singhai et al., 2024).
  • Tropical cyclone intensity: Storms feed on heat from the ocean surface. SST above 26.5 C significantly increases wind speed and rainfall of hurricanes / typhoons, causing significant damage to public safety and economy (Xu et al., 2016; Lavender et al., 2018).
@@
Sea level is monitored by tide gauges, such as by the NOAA Center [^3] https://tidesandcurrents.noaa.gov/ for Operational Oceanographic Products and Services (CO-OPS) in the United States, and globally via the University of Hawaii Sea Level Center [^4] https://uhslc.soest.hawaii.edu/, which provide access to decades of hourly and 6-minute water level observations.Distributed in CSV, NetCDF, or via API outputs (JSON/XML), these records are essential but require specialized technical skills and tools to process (Moftakhari et al., 2015).More recently, a high-resolution NOAA Coastal Ocean ReAnalysis (CORA) (Rose et al., 2024; Keeney et al., 2025) enables a broader spatial representation of coastal sea level, however these data are similarly distributed in NetCDF and are large, with the entire data record exceeding 40TB. The large data size further complicates accessibility by requiring familiarity with cloud data storage and access. For example,Sea level data are critical for understanding related natural hazards such as:
  • High tide flooding: Sea level rise has made ordinary high tides sufficient to flood low-lying urban areas (e.g., Boston, Norfolk), causing increasingly frequent “sunny-day” floods that disrupt traffic and infrastructure (Li et al., 2021; Hino et al., 2019).
  • Storm surge risk: Elevated baseline sea levels combined with coastal storms dramatically worsen flood impact during hurricanes and typhoons (Lavender et al., 2018).
  • Long-term rise: Tide gauge data indicate global sea levels have increased by about 17cm from 1920-2020, contributing to more frequent and severe coastal flooding events (Sweet et al., 2022).
@@ 3.1 Large Language Models (LLMs)
Large language models such as GPT-4 (OpenAI et al., 2024), LLaMA (Touvron et al., 2023), and PaLM (Chowdhery et al., 2022) are transformer-based neural networks (Vaswani et al., 2023) trained on large-scale text corpora to predict the next token in a sequence. This training paradigm enables them to generate coherent paragraphs, answer complex questions, and reason over natural language prompts. Despite their impressive fluency, pretrained LLMs rely solely on patterns learned during training and have no inherent access to real-time or verifiable data sources.As a result, they may produce plausible-sounding but inaccurate outputs—a limitation known as hallucination (Ji et al., 2023). This issue is especially problematic in scientific domains, where factual correctness and transparency are essential.
@@ 3.1.1 Embeddings and Semantic Representations
To support semantic reasoning and information retrieval, AI systems frequently use vector-based representations known as embeddings. An embedding is a numeric vector that encodes the meaning of a word, sentence, or document in a high-dimensional space. Similar concepts are mapped to geometrically close vectors, enabling systems to retrieve semantically related content even when surface forms differ. Early embedding models like Word2Vec (Mikolov et al., 2013) and GloVe (Pennington et al., 2014) assigned static vectors to words, whereas newer models such as BERT (Devlin et al., 2019) and Sentence-BERT (Reimers and Gurevych, 2019) generate contextual embeddings, allowing the same word to have different vector representations depending on its usage. These embeddings serve as an interface between human language and machine reasoning across tasks such as search, classification, and clustering.
@@ 3.2 Retrieval-Augmented Generation (RAG)
Retrieval-Augmented Generation (RAG) enhances the factuality of LLM outputs by integrating external document retrieval into the generation pipeline (Lewis et al., 2021). A RAG system typically follows a two-stage process:
  • The user query is embedded and compared against an external corpus to retrieve relevant documents (Karpukhin et al., 2020).
  • The retrieved content is passed into the LLM’s context window, allowing the model to generate responses grounded in real evidence (Izacard and Grave, 2021).
This architecture improves transparency and reduces hallucination by enabling responses conditioned on externally retrieved and verifiable information—making RAG particularly useful for scientific and high-stakes domains (Gao et al., 2024).
@@ 3.3 Agentic RAG and Tool Use
While classical RAG pipelines rely on retrieval alone, many real-world tasks—such as querying scientific datasets, generating visualizations, or running domain-specific simulations—require interaction with specialized computational tools. Agentic RAG extends the RAG paradigm by enabling models to plan, select, and invoke tools dynamically in response to user queries (Masterman et al., 2024). This design allows the system to chain multiple steps, combining retrieval, computation, and synthesis into a coherent workflow.
Figure 2: Function-calling workflow in large language models (LLMs). The LLM parses a user query (e.g., What is the sea level in Boston in 2021?) into a structured API call (e.g., get_ssh(2021, Boston)). The call is executed externally, and the returned results (e.g., figures or statistics) are combined into the final response (OpenAI, 2023).
Function Calling. A key mechanism enabling agentic behavior is function calling, where the model emits structured outputs (typically in JSON [^5] A lightweight, text-based format for structured data exchange. JSON (JavaScript Object Notation) is commonly used to encode hierarchical data such as parameter sets, metadata, and API calls, and is well suited for integrating machine learning models with scientific tools and datasets. format) to request downstream tool execution (Schick et al., 2023; OpenAI, 2023).As illustrated in Figure 2, this process follows a loop of interpretation, execution, and synthesis: a user query such as “What is the sea level in Boston in 2021?” is first translated by the model into a structured API call (e.g., get_ssh("2021", "Boston")). The external function is then executed—retrieving sea surface height (SSH) data—and the result (e.g., image links and numerical statistics) is returned to the model. The LLM integrates this structured output into a natural language response that is both accurate and verifiable. This approach enables LLMs to act as dynamic interfaces to computational infrastructures, making it possible to construct interactive and trustworthy AI systems for scientific and engineering domains. By incorporating real-time function execution and grounding model outputs in factual data, function calling enhances both the transparency and utility of LLM-driven workflows.
@@ 4 Related Work
LLMs in Scientific Domains. A central challenge in applying large language models (LLMs) to scientific domains lies in ensuring factual accuracy, reproducibility, and timeliness (Ji et al., 2023). Although LLMs excel in language fluency and general reasoning, their knowledge is static and often opaque, making it difficult to verify or update information. This limitation is critical for scientific applications, where outputs must be grounded in authoritative, up-to-date datasets and follow domain-specific standards. In climate and Earth sciences, for example, models that cannot handle structured formats such as NetCDF or GRIB risk misinterpretation of quantitative results. Recent work in biomedical (Singhal et al., 2023) and materials science (Mostafa et al., 2024) has similarly shown that domain adaptation is essential for reliable use in high-stakes research contexts.
Tool Augmentation and Retrieval-Augmented Generation. A widely studied mitigation strategy is tool augmentation, where LLMs invoke external tools for information retrieval or computation (Schick et al., 2023). Toolformer and related agent-based systems such as HuggingGPT (Shen et al., 2023) demonstrate that models can be trained to autonomously call APIs such as search engines, calculators, or code execution environments, improving factuality. Retrieval-augmented generation (RAG) further enhances this by grounding responses in retrieved documents, as seen in WebGPT (Nakano et al., 2022). However, these frameworks primarily operate on unstructured text and lack native support for specialized scientific data formats or spatiotemporal datasets.
LLMs for Earth and Ocean Sciences. In the Earth sciences, LLM integration is still nascent. Domain-specific models such as ClimateGPT (Thulke et al., 2024) and GeoGalactica (Lin et al., 2024) have shown promise in processing literature and reports, but they do not incorporate real-time observational data streams. In the ocean domain, OceanGPT (Bi et al., 2024) fine-tunes large language models on a multi-agent–generated instruction dataset and introduces OCEANBENCH for benchmarking 15 ocean science tasks. While it improves domain-specific reasoning compared with general LLMs, it operates primarily on static corpora and lacks real-time integration with authoritative observational datasets, limiting its applicability for time-sensitive reporting.A recent study further introduced the Intelligent Data Exploring Assistant (IDEA), which integrates LLMs with domain-specific geoscience data and analytical tools (Widlansky and Komar, 2025).

### v2

@@ 1 Introduction
The ocean plays a pivotal role in regulating Earth’s climate, supporting biodiversity, and sustaining coastal economies. Accurate understanding of oceanographic processes underpins climate forecasting, hazard preparedness, and sustainable resource management (hurrell2009unified; roemmich2019argo). Over the past two decades, advances in satellite remote sensing, in-situ observation networks (e.g., Argo (Wong2020-rk)), and reanalysis programs(e.g., ERA5 (ERA5) and GLORYS (GLORYS)) have dramatically expanded the availability of high-resolution ocean data. However, these datasets remain technically challenging to access and work with due to their complex formats and spatiotemporal structure, creating barriers for non-expert users and overhead even for domain experts.
Existing tools for accessing oceanographic information fall broadly into three categories:general-purpose large language models (LLMs), domain-specific data servers, and hybrid geospatial analysis environments. General-purpose large language models (LLMs), such as GPT-4 (openai2024gpt4technicalreport) and LLaMA (touvron2023llamaopenefficientfoundation), offer intuitive natural language interaction but lack direct access to authoritative datasets. Without integration to structured data sources, these models cannot reliably execute parameterized scientific queries, and their outputs are prone to hallucination ji2023survey, producing unverifiable or incorrect statements. Domain-specific platforms, such as the NOAA THREDDS Data Server [^1] https://www.ncei.noaa.gov/thredds/catalog/catalog.html and the NOAA Open Data Dissemination Program [^2] https://www.noaa.gov/information-technology/open-data-dissemination, provide structured, transparent access to observational and model data but require substantial technical expertise to navigate file hierarchies, parse specialized formats (e.g., NetCDF, GRIB), and perform downstream analysis. Hybrid geospatial analysis environments, such as Google Earth Engine (googleEarth), provide access to global geospatial datasets—including some oceanographic layers—and support large-scale data processing and interactive visualization. However, they typically require users to write JavaScript or Python code and do not offer conversational interfaces or direct parameterized querying of authoritative ocean datasets.
@@ 2 Ocean Background
Sea surface temperature (SST) is widely used in earth sciences from ocean ecosystem conservation to weather forecasting. SST is usually measured by satellite sensors as well as in-situ platforms such as buoys and drifters. Prominent sources include NOAA’s Coral Reef Watch and the Group for High Resolution Sea Surface Temperature(GHRSST), which offer daily global SST data at 5 km and finer resolution in NetCDF format. These files include spatial/temporal coordinates, quality flags, and uncertainty estimates, requiring coding expertise with scientific data packages such as xarray or netCDF4 for access (crw).
@@
  • El Niño-Southern Oscillation (ENSO): Characterized by prolonged SST anomalies in the eastern tropical Pacific, ENSO events influence global atmospheric circulation and provoke extreme weather like droughts in South Asia and flooding in South America (elNinoReview; turn0search0).
  • Monsoon variability: Warmer SSTs in the Indian and Pacific Oceans enhance monsoon rainfall; cooler SSTs can suppress or delay monsoon onset, hindering agricultural production and affecting large populations (turn0search0; SINGHAI2024).
  • Tropical cyclone intensity: Storms feed on heat from the ocean surface. SST above 26.5 C significantly increases wind speed and rainfall of hurricanes / typhoons, causing significant damage to public safety and economy (Xu2016_SST_TC; nhess-18-795-2018).
@@
Sea level is monitored by tide gauges, such as by the NOAA Center [^3] https://tidesandcurrents.noaa.gov/ for Operational Oceanographic Products and Services (CO-OPS) in the United States, and globally via the University of Hawaii Sea Level Center [^4] https://uhslc.soest.hawaii.edu/, which provide access to decades of hourly and 6-minute water level observations.Distributed in CSV, NetCDF, or via API outputs (JSON/XML), these records are essential but require specialized technical skills and tools to process (moftakhari2015nuisance).More recently, a high-resolution NOAA Coastal Ocean ReAnalysis (CORA) (CORA_data; CORA_report) enables a broader spatial representation of coastal sea level, however these data are similarly distributed in NetCDF and are large, with the entire data record exceeding 40TB. The large data size further complicates accessibility by requiring familiarity with cloud data storage and access. For example,Sea level data are critical for understanding related natural hazards such as:
  • High tide flooding: Sea level rise has made ordinary high tides sufficient to flood low-lying urban areas (e.g., Boston, Norfolk), causing increasingly frequent “sunny-day” floods that disrupt traffic and infrastructure (li2021evolving; Hino2019-zf).
  • Storm surge risk: Elevated baseline sea levels combined with coastal storms dramatically worsen flood impact during hurricanes and typhoons (nhess-18-795-2018).
  • Long-term rise: Tide gauge data indicate global sea levels have increased by about 17cm from 1920-2020, contributing to more frequent and severe coastal flooding events (Sweet2022SLR).
@@ 3.1 Large Language Models (LLMs)
Large language models such as GPT-4 (openai2024gpt4technicalreport), LLaMA (touvron2023llamaopenefficientfoundation), and PaLM (chowdhery2022palm) are transformer-based neural networks (vaswani2023attentionneed) trained on large-scale text corpora to predict the next token in a sequence. This training paradigm enables them to generate coherent paragraphs, answer complex questions, and reason over natural language prompts. Despite their impressive fluency, pretrained LLMs rely solely on patterns learned during training and have no inherent access to real-time or verifiable data sources.As a result, they may produce plausible-sounding but inaccurate outputs—a limitation known as hallucination (ji2023survey). This issue is especially problematic in scientific domains, where factual correctness and transparency are essential.
@@ 3.1.1 Embeddings and Semantic Representations
To support semantic reasoning and information retrieval, AI systems frequently use vector-based representations known as embeddings. An embedding is a numeric vector that encodes the meaning of a word, sentence, or document in a high-dimensional space. Similar concepts are mapped to geometrically close vectors, enabling systems to retrieve semantically related content even when surface forms differ. Early embedding models like Word2Vec (mikolov2013efficient) and GloVe (pennington2014glove) assigned static vectors to words, whereas newer models such as BERT (devlin2019bert) and Sentence-BERT (reimers2019sentence) generate contextual embeddings, allowing the same word to have different vector representations depending on its usage. These embeddings serve as an interface between human language and machine reasoning across tasks such as search, classification, and clustering.
@@ 3.2 Retrieval-Augmented Generation (RAG)
Retrieval-Augmented Generation (RAG) enhances the factuality of LLM outputs by integrating external document retrieval into the generation pipeline (lewis2020retrieval). A RAG system typically follows a two-stage process:
  • The user query is embedded and compared against an external corpus to retrieve relevant documents (karpukhin2020dense).
  • The retrieved content is passed into the LLM’s context window, allowing the model to generate responses grounded in real evidence (izacard2021leveraging).
This architecture improves transparency and reduces hallucination by enabling responses conditioned on externally retrieved and verifiable information—making RAG particularly useful for scientific and high-stakes domains (gao2024retrieval).
@@ 3.3 Agentic RAG and Tool Use
While classical RAG pipelines rely on retrieval alone, many real-world tasks—such as querying scientific datasets, generating visualizations, or running domain-specific simulations—require interaction with specialized computational tools. Agentic RAG extends the RAG paradigm by enabling models to plan, select, and invoke tools dynamically in response to user queries (masterman2024landscapeemergingaiagent). This design allows the system to chain multiple steps, combining retrieval, computation, and synthesis into a coherent workflow.
Figure 2: Function-calling workflow in large language models (LLMs). The LLM parses a user query (e.g., What is the sea level in Boston in 2021?) into a structured API call (e.g., get_ssh(2021, Boston)). The call is executed externally, and the returned results (e.g., figures or statistics) are combined into the final response (openai_function_call).
Function Calling. A key mechanism enabling agentic behavior is function calling, where the model emits structured outputs (typically in JSON [^5] A lightweight, text-based format for structured data exchange. JSON (JavaScript Object Notation) is commonly used to encode hierarchical data such as parameter sets, metadata, and API calls, and is well suited for integrating machine learning models with scientific tools and datasets. format) to request downstream tool execution (schick2023toolformer; openai_function_call).As illustrated in Figure 2, this process follows a loop of interpretation, execution, and synthesis: a user query such as “What is the sea level in Boston in 2021?” is first translated by the model into a structured API call (e.g., get_ssh("2021", "Boston")). The external function is then executed—retrieving sea surface height (SSH) data—and the result (e.g., image links and numerical statistics) is returned to the model. The LLM integrates this structured output into a natural language response that is both accurate and verifiable. This approach enables LLMs to act as dynamic interfaces to computational infrastructures, making it possible to construct interactive and trustworthy AI systems for scientific and engineering domains. By incorporating real-time function execution and grounding model outputs in factual data, function calling enhances both the transparency and utility of LLM-driven workflows.
@@ 4 Related Work
LLMs in Scientific Domains. A central challenge in applying large language models (LLMs) to scientific domains lies in ensuring factual accuracy, reproducibility, and timeliness (ji2023survey). Although LLMs excel in language fluency and general reasoning, their knowledge is static and often opaque, making it difficult to verify or update information. This limitation is critical for scientific applications, where outputs must be grounded in authoritative, up-to-date datasets and follow domain-specific standards. In climate and Earth sciences, for example, models that cannot handle structured formats such as NetCDF or GRIB risk misinterpretation of quantitative results. Recent work in biomedical (singhal2023large) and materials science (mostafa2024grag) has similarly shown that domain adaptation is essential for reliable use in high-stakes research contexts.
Tool Augmentation and Retrieval-Augmented Generation. A widely studied mitigation strategy is tool augmentation, where LLMs invoke external tools for information retrieval or computation (schick2023toolformer). Toolformer and related agent-based systems such as HuggingGPT (shen2023hugginggpt) demonstrate that models can be trained to autonomously call APIs such as search engines, calculators, or code execution environments, improving factuality. Retrieval-augmented generation (RAG) further enhances this by grounding responses in retrieved documents, as seen in WebGPT (nakano2022webgpt). However, these frameworks primarily operate on unstructured text and lack native support for specialized scientific data formats or spatiotemporal datasets.
LLMs for Earth and Ocean Sciences. In the Earth sciences, LLM integration is still nascent. Domain-specific models such as ClimateGPT (thulke2024climategpt) and GeoGalactica (lin2024geoga) have shown promise in processing literature and reports, but they do not incorporate real-time observational data streams. In the ocean domain, OceanGPT (bi2024oceangpt) fine-tunes large language models on a multi-agent–generated instruction dataset and introduces OCEANBENCH for benchmarking 15 ocean science tasks. While it improves domain-specific reasoning compared with general LLMs, it operates primarily on static corpora and lacks real-time integration with authoritative observational datasets, limiting its applicability for time-sensitive reporting.A recent study further introduced the Intelligent Data Exploring Assistant (IDEA), which integrates LLMs with domain-specific geoscience data and analytical tools (Widlansky2025).